OpenStudy (anonymous):

Why ln(i)=(1/2)*pi*i?

OpenStudy (anonymous):

The situation actually was \[i^i=e^{i*lni}=e^{i(1/2)\pi*i}=e^{-(1/2)\pi}\]

OpenStudy (earthcitizen):

yh, use Euler's formula

OpenStudy (earthcitizen):

\[e ^{i}e ^{lni}\]

OpenStudy (anonymous):

How does \[e^{i \pi}=-1\] help here, or is there another part of the formula that I;m missing?

OpenStudy (earthcitizen):

\[lni=i \pi/2\]

OpenStudy (anonymous):

OK, Thanks. Could you point me to an article with the proof?

OpenStudy (earthcitizen):

http://www.wolframalpha.com/input/?i=lni